OpenStudy (anonymous):

Rhode Island is the smalles state in the United States. Its land area is about 2.9x10^10 square feet. Alaska, the largest state, is about 5.5x10^2 times as large as Rhode Island. What is the land area of Alaska in square feet? Write your answer in scientific notation

OpenStudy (bahrom7893):

5.5*10^2*2.9*10^10

OpenStudy (bahrom7893):

= 15.95 * 10^12 = 1.6 * 10^13

OpenStudy (campbell_st):

multiply the 2 numbers ( 2.9x10^10)x(5.5 x 10^2) looks like 1.15 x 10^13 at a rough guess.